OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1/2x − 2, find f[g(−6)]. 9 3 −9 −3

OpenStudy (ghazi):

first find f(g(x)) put x=(1/2) x-2 in f(x) f(g(x))= (1/2)x-2-4= (1/2)x-6 now to find f(g(-6)) put x=-6 ... f(g(-6))= (1/2)* (-6)-6= -3-6=-9